# A Unified Framework for Inference with General Missingness Patterns and Machine Learning Imputation

### v1

@@ 5 Conclusion and Discussion
On the other hand, we know that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{WCC}}-\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}(\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k})=\widehat{{\bm{c}}}^{\top}\widehat{\bm{\vartheta}}$, where $\widehat{{\bm{c}}}=(\mathbf{I},-\widehat{\mathbf{W}}_{1},\cdots,-\widehat{\mathbf{W}}_{K},\widehat{\mathbf{W}}_{1},\cdots,\widehat{\mathbf{W}}_{K},\bm{0})^{\top}$. Since $\widehat{\mathbf{W}}_{k}\overset{p}{\rightarrow}\mathbf{W}_{k}$, we have $\widehat{{\bm{c}}}\overset{p}{\rightarrow}{\bm{c}}$, where ${\bm{c}}=(\mathbf{I},-\mathbf{W}_{1},\cdots,-\mathbf{W}_{K},\mathbf{W}_{1},\cdots,\mathbf{W}_{K},\bm{0})^{\top}$. Based on Proposition 1, we have both ${\bm{\theta}}^{\ast}=\widehat{{\bm{c}}}^{\top}\bm{\vartheta}^{\ast}$ and ${\bm{\theta}}^{\ast}={\bm{c}}^{\top}\bm{\vartheta}^{\ast}$. Therefore, we can obtain the asymptotic linear expansion of $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$:
$\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})=\widehat{{\bm{c}}}^{\top}\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-{\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{(i)}+o_{p}(1),$
where
${\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}=\begin{bmatrix}\dot{{\bm{\psi}}}^{-1}\\[6.0pt] -\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\\[6.0pt] \vdots\\ -\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{K}\\[6.0pt] \mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\\[6.0pt] \cdots\\ \mathbf{W}_{K}\dot{{\bm{\psi}}}^{-1}_{K}\\[6.0pt] \left(\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}\end{bmatrix}^{\top}$
@@
As a linear transformation of $\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})$, we know that $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ is asymptotically normal. We now derive the form of its asymptotic variance. Denote ${\bm{\psi}}^{(i)}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})$, and ${\bm{\psi}}^{(i)}_{k}={\bm{\psi}}(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}})$; since $\widehat{\mathbf{W}}_{k}\overset{p}{\rightarrow}\mathbf{W}_{k}$, we have:
@@
$=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left[-\dot{{\bm{\psi}}}^{-1}{\bm{\psi}}^{(i)}_{\text{WCC}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}({\bm{\psi}}^{(i)}_{1,k}-{\bm{\psi}}^{(i)}_{2,k})\right$.
$\;\;\;+\left.\left(\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}{\bm{m}}^{(i)}\right]+o_{p}(1)$
$=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left[\underbrace{-\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{(i)}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\widehat{\pi}_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\widehat{\pi}_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}^{(i)}_{k}}_{{\bm{\Gamma}}_{i}}\right$.
$\;\;\;+\left.\left(\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}{\bm{m}}^{(i)}\right]+o_{p}(1).$
Notice that,
$\dot{{\bm{\Gamma}}}_{\bm{\omega}}:=\mathbb{E}\left[\frac{\partial{\bm{\Gamma}}_{i}}{\partial{\bm{\omega}}}\Bigr{|}_{{\bm{\omega}}={\bm{\omega}}^{\ast}}\right]=\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k}).$
As a result,
$\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left({\bm{\Gamma}}_{i}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}_{i}\right)+o_{p}(1),$
@@
$\varphi_{i}={\bm{\Gamma}}_{i}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}_{i}.$
@@
$\text{Var}(\varphi_{i})=\mathbb{E}\left[\left({\bm{\Gamma}}_{i}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}_{i}\right)\left({\bm{\Gamma}}_{i}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}_{i}\right)^{\top}\right]$ (S1.1)
@@
${\bm{\Gamma}}_{i}=-\underbrace{\dot{{\bm{\psi}}}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{(i)}}_{\mathbf{A}}+\underbrace{\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}^{(i)}_{k}}_{\mathbf{B}}.$
Correspondingly, the asymptotic variance becomes $\text{Var}({\bm{\Gamma}}_{i})=\mathbb{E}\left[{\bm{\Gamma}}_{i}{\bm{\Gamma}}^{\top}_{i}\right]=\mathbb{E}[\mathbf{A}\mathbf{A}^{\top}-\mathbf{A}\mathbf{B}^{\top}-\mathbf{B}\mathbf{A}^{\top}+\mathbf{B}\mathbf{B}^{\top}]$. Here,
$\mathbf{A}\mathbf{A}^{\top}$ $=\dot{{\bm{\psi}}}^{-1}{\bm{\psi}}{\bm{\psi}}^{\top}\dot{{\bm{\psi}}}^{-1\top}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})},$
$\mathbf{A}\mathbf{B}^{\top}$ $=\sum^{K}_{k=1}\dot{{\bm{\psi}}}^{-1}{\bm{\psi}}{\bm{\psi}}^{\top}_{k}\dot{{\bm{\psi}}}^{-1\top}_{k}\mathbf{W}^{\top}_{k}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right)$
$=\sum^{K}_{k=1}\dot{{\bm{\psi}}}^{-1}{\bm{\psi}}{\bm{\psi}}^{\top}_{k}\dot{{\bm{\psi}}}^{-1\top}_{k}\mathbf{W}^{\top}_{k}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})},$
$\mathbf{B}\mathbf{B}^{\top}$ $=\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}{\bm{\psi}}_{k}{\bm{\psi}}^{\top}_{k^{\prime}}\dot{{\bm{\psi}}}^{-1\top}_{k^{\prime}}\mathbf{W}^{\top}_{k^{\prime}}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})}+\frac{I({\mathcal{C}}_{i}=k)I({\mathcal{C}}_{i}=k^{\prime})}{\pi_{k}(G_{k}(\bm{Y}_{i}))\pi_{k^{\prime}}(G_{k^{\prime}}(\bm{Y}_{i}))}\right)$
@@
$\mathbb{E}\left[\mathbf{A}\mathbf{A}^{\top}\right]$ $=\underbrace{\dot{{\bm{\psi}}}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}}_{\Sigma_{{\bm{\theta}}}},$ (S1.2)
$\mathbb{E}\left[\mathbf{A}\mathbf{B}^{\top}\right]$ $=\sum^{K}_{k=1}\underbrace{\dot{{\bm{\psi}}}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}_{k}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{k}}_{\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}}\mathbf{W}^{\top}_{k},$ (S1.3)
$\mathbb{E}\left[\mathbf{B}\mathbf{B}^{\top}\right]$ $=\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\underbrace{\dot{{\bm{\psi}}}^{-1}_{k}\mathbb{E}\left[\frac{{\bm{\psi}}_{k}{\bm{\psi}}^{\top}_{k^{\prime}}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{k^{\prime}}}_{\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}}\mathbf{W}^{\top}_{k^{\prime}}$ (S1.4)
$+\sum^{K}_{k=1}\mathbf{W}_{k}\underbrace{\dot{{\bm{\psi}}}^{-1}_{k}\mathbb{E}\left[\frac{{\bm{\psi}}_{k}{\bm{\psi}}^{\top}_{k}}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right]\dot{{\bm{\psi}}}^{-1\top}_{k}}_{\Sigma_{{\bm{\gamma}}_{2,k}}}\mathbf{W}^{\top}_{k}.$ (S1.5)
@@
S1.4 Instruction of estimating $\Sigma_{\text{PS-PPI}}$.
To obtain a consistent estimate of $\Sigma_{\text{PS-PPI}}$, we may want to obtain consistent estimates for both the meat (e.g., $\mathbb{E}\left[\frac{{\bm{\psi}}{\bm{\psi}}^{\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]$ for $\Sigma_{\bm{\theta}}$) and the bread (e.g., $\dot{{\bm{\psi}}}^{-1}$ and $\dot{{\bm{\psi}}}^{-1\top}$ for $\Sigma_{\bm{\theta}}$) parts for each of the variance-covariance matrices $\Sigma_{{\bm{\theta}}}$, $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\Sigma_{{\bm{\gamma}}_{2,k}}$. A straightforward approach is to use the sample average over the fully-observed data to estimate the population average shown in both the meat and the bread parts. This can be done once the closed-form of $\bm{\psi}$${\bm{\psi}}_{k}$ and $\dot{{\bm{\psi}}}$$\dot{{\bm{\psi}}}_{k}$ can be derived. In the following we provide two examples, linear regression and logistic regression, for reference.

### v2

@@ Acknowledgment
From van2000asymptotic and Assumption 2, $\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})$ is asymptotically normal. To show that $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ is asymptotically normal, we want to show that it is $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is a linear combination of the elements in $\widehat{\bm{\vartheta}}$. We know that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{WCC}}-\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}(\widehat{{\bm{\gamma}}}_{1,k}-\widehat{{\bm{\gamma}}}_{2,k})=\widehat{{\bm{c}}}^{\top}\widehat{\bm{\vartheta}}$, where $\widehat{{\bm{c}}}=(\mathbf{I},-\widehat{\mathbf{W}}_{1},\cdots,-\widehat{\mathbf{W}}_{K},\widehat{\mathbf{W}}_{1},\cdots,\widehat{\mathbf{W}}_{K},\bm{0})^{\top}$. On the other hand, since the true parameter of interest $\bm{\theta}$$^{\ast}$ is the first element in $\bm{\vartheta}^{\ast}$, we can rewrite it as ${\bm{\theta}}^{\ast}={\bm{\theta}}^{\ast}-\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}{\bm{\gamma}}^{\ast}_{k}+\sum^{K}_{k=1}\widehat{\mathbf{W}}_{k}{\bm{\gamma}}^{\ast}_{k}=\widehat{{\bm{c}}}^{\top}\bm{\vartheta}^{\ast}$. Define ${\bm{c}}=(\mathbf{I},-\mathbf{W}_{1},\cdots,-\mathbf{W}_{K},\mathbf{W}_{1},\cdots,\\ \mathbf{W}_{K},\bm{0})^{\top}$; since $\widehat{\mathbf{W}}_{k}\overset{p}{\rightarrow}\mathbf{W}_{k}$, we have $\widehat{{\bm{c}}}\overset{p}{\rightarrow}{\bm{c}}$. The asymptotic linear expansion of $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ becomes:
$\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})=\widehat{{\bm{c}}}^{\top}\sqrt{N}(\widehat{\bm{\vartheta}}-\bm{\vartheta}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-{\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{\ast(i)}+o_{p}(1).$
@@
Hence, $\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})$ is asymptotically normal. We now derive the form of its asymptotic variance. First,
${\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}=\begin{bmatrix}\dot{{\bm{\psi}}}_{\infty}^{-1}\\[6.0pt] -\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\\[6.0pt] \vdots\\ -\mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{K}\\[6.0pt] \mathbf{W}_{1}\dot{{\bm{\psi}}}^{-1}_{1}\\[6.0pt] \cdots\\ \mathbf{W}_{K}\dot{{\bm{\psi}}}^{-1}_{K}\\[6.0pt] \left(\dot{{\bm{\psi}}}_{\infty}^{-1}\dot{{\bm{\psi}}}_{\infty}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}\end{bmatrix}^{\top}$
Denote ${\bm{\psi}}^{(i)}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}})$, and ${\bm{\psi}}^{(i)}_{k}={\bm{\psi}}(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}})$; similarly we have ${\bm{\psi}}^{\ast(i)}={\bm{\psi}}(\bm{Y}_{i};{\bm{\theta}}^{\ast})$ and ${\bm{\psi}}^{\ast(i)}_{k}={\bm{\psi}}(\bm{Y}^{(k)}_{\text{imp},i};{\bm{\gamma}}_{k}^{\ast})$. Since $\widehat{\mathbf{W}}_{k}\overset{p}{\rightarrow}\mathbf{W}_{k}$, we have:
@@
$=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}-{\bm{c}}^{\top}\dot{\bm{\Psi}}^{-1}\bm{\Psi}^{\ast(i)}+o_{p}(1)$
$=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left[-\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}_{\text{WCC}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}({\bm{\psi}}^{\ast(i)}_{1,k}-{\bm{\psi}}^{\ast(i)}_{2,k})\right$.
$\;\;\;+\left.\left(\dot{{\bm{\psi}}}_{\infty}^{-1}\dot{{\bm{\psi}}}_{\infty}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k})\right)\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}\right]+o_{p}(1).$ (S1.1)
Denote ${\bm{\Gamma}}^{(i)}=-\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{(i)}_{\text{WCC}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}({\bm{\psi}}^{(i)}_{1,k}-{\bm{\psi}}^{(i)}_{2,k})$. Note that this is a function of $\bm{\vartheta}$. Then we have
$\dot{{\bm{\Gamma}}}_{\bm{\omega}}:=\mathbb{E}\left[\frac{\partial{\bm{\Gamma}}^{(i)}}{\partial{\bm{\omega}}}\Bigr|_{\bm{\vartheta}=\bm{\vartheta}^{\ast}}\right]=\dot{{\bm{\psi}}}^{-1}\dot{{\bm{\psi}}}^{\bm{\omega}}-\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}(\dot{{\bm{\psi}}}^{\bm{\omega}}_{1,k}-\dot{{\bm{\psi}}}^{\bm{\omega}}_{2,k}).$
As a result, from Eq. (S1.1) we can obtain:
$\sqrt{N}(\widehat{{\bm{\theta}}}_{\text{PS-PPI}}-{\bm{\theta}}^{\ast})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\left({\bm{\Gamma}}^{\ast(i)}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}\right)+o_{p}(1),$
@@
$\varphi_{i}={\bm{\Gamma}}^{\ast(i)}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}.$
@@
$\text{Var}(\varphi_{i})=\mathbb{E}\left[\left({\bm{\Gamma}}^{\ast(i)}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}\right)\left({\bm{\Gamma}}^{\ast(i)}+\dot{{\bm{\Gamma}}}_{\bm{\omega}}\mathbf{M}^{-1}{\bm{m}}^{\ast(i)}\right)^{\top}\right]$ (S1.2)
@@
${\bm{\Gamma}}^{\ast(i)}$ $=-\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}_{\text{WCC}}+\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}({\bm{\psi}}^{\ast(i)}_{1,k}-{\bm{\psi}}^{\ast(i)}_{2,k})$
$=-\underbrace{\dot{{\bm{\psi}}}_{\infty}^{-1}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}{\bm{\psi}}^{\ast(i)}}_{\mathbf{A}^{(i)}}+\underbrace{\sum^{K}_{k=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right){\bm{\psi}}^{\ast(i)}_{k}}_{\mathbf{B}^{(i)}}.$
Correspondingly, the asymptotic variance becomes $\text{Var}({\bm{\Gamma}}^{\ast(i)})=\mathbb{E}\left[{\bm{\Gamma}}^{\ast(i)}{\bm{\Gamma}}^{\ast(i)\top}\right]=\mathbb{E}[\mathbf{A}^{(i)}\mathbf{A}^{(i)\top}-\mathbf{A}^{(i)}\mathbf{B}^{(i)\top}-\mathbf{B}^{(i)}\mathbf{A}^{(i)\top}+\mathbf{B}^{(i)}\mathbf{B}^{(i)\top}]$. Here,
$\mathbf{A}^{(i)}\mathbf{A}^{(i)\top}$ $=\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}\dot{{\bm{\psi}}}_{\infty}^{-1\top}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})},$
$\mathbf{A}^{(i)}\mathbf{B}^{(i)\top}$ $=\sum^{K}_{k=1}\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}_{k}\dot{{\bm{\psi}}}^{-1\top}_{k}\mathbf{W}^{\top}_{k}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi_{\infty}(\bm{Y}_{i})}-\frac{I({\mathcal{C}}_{i}=k)}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right)$
$=\sum^{K}_{k=1}\dot{{\bm{\psi}}}_{\infty}^{-1}{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}_{k}\dot{{\bm{\psi}}}^{-1\top}_{k}\mathbf{W}^{\top}_{k}\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})},$
$\mathbf{B}^{(i)}\mathbf{B}^{(i)\top}$ $=\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\dot{{\bm{\psi}}}^{-1}_{k}{\bm{\psi}}^{\ast(i)}_{k}{\bm{\psi}}^{\ast(i)\top}_{k^{\prime}}\dot{{\bm{\psi}}}^{-1\top}_{k^{\prime}}\mathbf{W}^{\top}_{k^{\prime}}\left(\frac{I({\mathcal{C}}_{i}=\infty)}{\pi^{2}_{\infty}(\bm{Y}_{i})}+\frac{I({\mathcal{C}}_{i}=k)I({\mathcal{C}}_{i}=k^{\prime})}{\pi_{k}(G_{k}(\bm{Y}_{i}))\pi_{k^{\prime}}(G_{k^{\prime}}(\bm{Y}_{i}))}\right)$
@@
$\mathbb{E}\left[\mathbf{A}^{(i)}\mathbf{A}^{(i)\top}\right]$ $=\underbrace{\dot{{\bm{\psi}}}_{\infty}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}_{\infty}^{-1\top}}_{\Sigma_{{\bm{\theta}}}},$ (S1.3)
$\mathbb{E}\left[\mathbf{A}^{(i)}\mathbf{B}^{(i)\top}\right]$ $=\sum^{K}_{k=1}\underbrace{\dot{{\bm{\psi}}}_{\infty}^{-1}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}_{k}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{k}}_{\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}}\mathbf{W}^{\top}_{k},$ (S1.4)
$\mathbb{E}\left[\mathbf{B}^{(i)}\mathbf{B}^{(i)\top}\right]$ $=\sum^{K}_{k=1}\sum^{K}_{k^{\prime}=1}\mathbf{W}_{k}\underbrace{\dot{{\bm{\psi}}}^{-1}_{k}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}_{k}{\bm{\psi}}^{\ast(i)\top}_{k^{\prime}}}{\pi_{\infty}(\bm{Y}_{i})}\right]\dot{{\bm{\psi}}}^{-1\top}_{k^{\prime}}}_{\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}}\mathbf{W}^{\top}_{k^{\prime}}$ (S1.5)
$+\sum^{K}_{k=1}\mathbf{W}_{k}\underbrace{\dot{{\bm{\psi}}}^{-1}_{k}\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}_{k}{\bm{\psi}}^{\ast(i)\top}_{k}}{\pi_{k}(G_{k}(\bm{Y}_{i}))}\right]\dot{{\bm{\psi}}}^{-1\top}_{k}}_{\Sigma_{{\bm{\gamma}}_{2,k}}}\mathbf{W}^{\top}_{k}.$ (S1.6)
@@
S1.4 Details of estimating $\Sigma_{\text{PS-PPI}}$.
To obtain a consistent estimate of $\Sigma_{\text{PS-PPI}}$, we may want to obtain consistent estimates for both the meat (e.g., $\mathbb{E}\left[\frac{{\bm{\psi}}^{\ast(i)}{\bm{\psi}}^{\ast(i)\top}}{\pi_{\infty}(\bm{Y}_{i})}\right]$ for $\Sigma_{\bm{\theta}}$) and the bread (e.g., $\dot{{\bm{\psi}}}_{\infty}^{-1}$ and $\dot{{\bm{\psi}}}_{\infty}^{-1\top}$ for $\Sigma_{\bm{\theta}}$) parts for each of the variance-covariance matrices $\Sigma_{{\bm{\theta}}}$, $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\Sigma_{{\bm{\gamma}}_{2,k}}$. This can be done once the closed-form expression of ${\bm{\psi}}^{\ast(i)}$, $\dot{{\bm{\psi}}}_{\infty}$, ${\bm{\psi}}^{\ast(i)}_{k}$, and $\dot{{\bm{\psi}}}_{k}$ can be derived. In the following, we provide two examples, linear regression and logistic regression, to illustrate the derivation of the closed-form of $\Sigma_{{\bm{\theta}}}$ from Eq. (S1.3). Similar derivations can be done for $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\Sigma_{{\bm{\gamma}}_{2,k}}$ when referring to Eq. (S1.4), (S1.5), and (S1.6).